OpenStudy (anonymous):

Solve for the variable. Reduce the answer to lowest terms. (5/8)h=160

OpenStudy (ash2326):

We have \[\frac 5 8 h=160\] multiply both sides by 8/5 \[\frac 58 \times \frac 85 h=160\times \frac 85\] \[\frac {\cancel 5}{\cancel8} \times \frac {\cancel8}{\cancel5} h=160\times \frac 85\] we get \[h=160\times \frac 85\] just simplify the right side and you'll get your answer